OpenStudy (geekfromthefutur):

A system of linear inequalities is shown below: y ≥ -2x - 2 y + x ≥ 0 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (geekfromthefutur):

@FlvsGirl

OpenStudy (anonymous):

@texaschic101 I gtg, so sorry....

OpenStudy (texaschic101):

y >= -2x - 2 slope is -2 y intercept is (0,-2) x intercept can be found by subbing in 0 for y.. (you do not need the inequality sign right now, so change it to just an equal sign 0 = -2x - 2 2x = -2 x = -1....so the x intercept is (-1,0) plot your intercepts...(0,-2) and (-1,0) so start at (-1,0) and since your slope is -2, go down 2 and to the right 1, then down 2, and to the right 1, and you should cross the y axis at (0,-2) it will be a solid line because we have an equal sign. It will be shaded above the line because it is greater then. now for the next line y + x >= 0 y >= -x + 0 slope is -1 y intercept is (0,0) x intercept is : y = -x + 0 -1x = 0 x = 0...so the x intercept is (0,0) so start at (0,0)...and since the slope is -1, go down 1 and to the right 1, and down 1 and to the right 1. It will be a solid line because we have the equal sign. It will be shaded above the line because it is greater then.

OpenStudy (geekfromthefutur):

oh ok wow thats alot is tht the answer ???..

OpenStudy (texaschic101):

That is the steps to graph it...

OpenStudy (geekfromthefutur):

uh... lol how exactly will i be able sort of put this in a little steps jw cuse thts alot

OpenStudy (texaschic101):

does it have to be for this exact problem, or can it be the steps to solve for any system of inequalities ?

OpenStudy (geekfromthefutur):

it has to be for tht pacific equation i wish i could any other system

OpenStudy (texaschic101):

lets make it shorter then... y >= -2x + 2 slope = -2 y intercept = (0,2) x intercept = (-1,0) start at (-1,0) and since slope is -2, go down 2 and to the right 1, and down 2 and to the right 1 and you should cross the y axis at (0,-2) solid line because it has an equal sign. Shading above the line because it is > then. Next line.. y + x >= 0 slope = -1 y intercept = (0,0) x intercept = (0,0) start at (0,0) and since slope is -1, go down 1, and to the right 1, then down 1 and to the right 1. solid line because it has an equal sign. Shading above the line because it is > then. That is the shortest I can make it

OpenStudy (texaschic101):

I am sorry but I have to go now....I hope I helped

OpenStudy (geekfromthefutur):

oh ok thts actually perfect :D ok so is tht for part a. ok yea u helped a lot tysm :D

OpenStudy (texaschic101):

:)

OpenStudy (geekfromthefutur):

:) ignore the *is tht part a* thing lol i was looking at a diff question lol but ty again :)

OpenStudy (geekfromthefutur):

oh hey btw can u delete the thing u posted js cuse just in case my teacher or a mod comes and sees it